Find the final amount of money in an account if $5,000 is deposited at 6% interest compounded weekly and the money is left for 7 years.
Topic - Interest- N Compound a Year

Answer by lwsshak3(11628)

Find the final amount of money in an account if $5,000 is deposited at 6% interest compounded weekly and the money is left for 7 years.
***
Compound Interest Formula: A=P(1+r/n)^nt, P=initial investment, r=interest rate, n=number of compounding periods per year, A=amt after t-years.
For given problem:
P=5000
r=.06
n=52
t=7
A=5000(1+.06/52)^(7*52)
=5000(1+0.001154)^364
=5000(1.001154)^364
≈7608
final amount of money in the account≈$7,608